OpenStudy (anonymous):

A boy of mass 50.7 kg is initially on a skateboard of mass 2.00 kg, moving at a speed of 10.4 m/s. The boy falls off the skateboard, and his center of mass moves forward at a speed of 11.1 m/s. Find the final velocity of the skateboard.

OpenStudy (jamesj):

Use Conservation of Momentum. Find the momentum of both boy and skateboard before and after the fall.

OpenStudy (anonymous):

could you show this because I not the one in the class my friend is and I'm just posting the question on here for him

OpenStudy (jamesj):

Your friend should be here then. I can explain it to him in detail.

OpenStudy (anonymous):

The equation for the conservation of momentum is this: m_1*v_1i + m_2*v_2i = m_1*v_1f + m_2*v_2f Where m = mass v = velocity i = initial f = final

OpenStudy (anonymous):

Applied to your question: (50.7)(10.4) + (2)(10.4) = (50.7)(11.1) + (2)v_2f (527.28) + (20.8) = (562.77) + (2)v_2f 548.08 = 562.77 + (2)v_2f -14.69 = (2)v_2f -7.345 = v_2f So the skateboard moves 7.345 m/s in the opposite direction.